# AQM: A Refresh of the Abstract Qubit Model for Quantum Co-design

### v1

@@ IV.2.2 Nitrogen-Vacancy centers
Nitrogen-vacancy centers Wrachtrup and Jelezko (2006); Doherty et al. (2011, 2013) combined with other solid-state defect centers have emerged as promising candidates for QIP. Defect color centers inside the solid state systems can be nicely fabricated and implanted inside the solid crystal, while they can have long coherence spin states that can be manipulated using microwave and optically readout.
Qubit states.
@@
The negatively charged NV centers have six electrons localized around the defect. The electronic ground state manifold consists of three spin-1 states. When there is no magnetic field applied, the state with $S_{1}=0$ and two states with $S_{z}=\pm 1$ splits by $2.87$ GHz Doherty et al. (2011, 2013). When the NV center electronic spin states are utilized to encode quantum information, a magnetic field is applied to further break the degeneracy of the two states with $S_{z}=\pm 1$, and use one of them with the $S_{z}=0$ state to form a qubit space Fuchs et al. (2009, 2010); Abobeih et al. (2018). Another way to encode quantum information is to encode into the nuclear spin states of the nearby atoms (Carbon or Nitrogen atoms) Maurer et al. (2012); Bradley et al. (2019); Pompili et al. (2021); Bartling et al. (2022); Hermans et al. (2022); Xie et al. (2023).
Operations.
@@
The universal control of a qubit encoded into NV center electronic states can be achieved by coherent microwave drives Fuchs et al. (2009, 2010). The universal control on the nuclear spin states can be achieved by addressing the nuclear spin states by manipulating the electronic states of the NV center via the hyperfine interaction Jelezko et al. (2004); Abobeih et al. (2022).
@@
However, implementing entangling gates between NV centers to support a full universal gate set for QC is challenging. Entanglement between two electronic spin qubits of NV centers can be entangled by photon heralded entanglement generation process Bernien et al. (2013); Pfaff et al. (2014); Hensen et al. (2015); Pompili et al. (2021). Entanglement gates between the electronic spin qubit and the nearby nuclear spin qubits can be performed by microwave drive on electronic spin degree of freedom Cramer et al. (2016); Taminiau et al. (2014, 2012); Dong et al. (2020); Abobeih et al. (2022).
@@
With the heralded entanglement between the electronic states of two NV centers, the entangled state can be used to enable gate operations between the nearby nuclear spin states. It can potentially be used to obtain a remote entangling gate for nuclear spin qubits, which can be used to build a quantum network Pompili et al. (2021).
Connectivity.
@@
Despite the possibility of heralded entanglement generation, two-qubit entanglement occurs relatively slowly, with a reported rate of $9$ Hz for remote entanglement between two NV centers Pompili et al. (2021). Although improvements in photon emission, collection, and detection efficiency could enhance generation speed, the speed of generating entanglement of NV qubits remains slower compared to trapped ion and superconducting systems.
Long coherence time.
@@
The electronic qubit of NV centers can preserve a long lifetime. Even at room temperature, the NV electronic states can preserve $1.8$ ms coherence time in an isotopically pure diamond crystal Balasubramanian et al. (2009). The nuclear spin states can have even longer lifetime Maurer et al. (2012); Bradley et al. (2019); Pompili et al. (2021); Bartling et al. (2022); Hermans et al. (2022); Xie et al. (2023). One challenge of NV center electronic states and nuclear spin states is the relatively short coherence time, which is due to the spin bath of the surrounding nuclear spins in the diamond crystal. However, the effect of the spin bath can be mitigated by dynamical decoupling Maurer et al. (2012); Bar-Gill et al. (2013); Abobeih et al. (2018, 2022); Bradley et al. (2019). With the help of dynamical decoupling, the coherence time of the electronic spin state can reach $1.58$ s Abobeih et al. (2018), while the nuclear spin states can be extended to $63$ s Bradley et al. (2019); Abobeih et al. (2022).
Qubit Readout.
@@
The measurement of NV center electronic states is implemented by optically pumping the states to the excited state manifold, and detecting the emitted photons. Due to the existence of a non-radiative relaxation path from the excited spin 00 states, the state of the NV electronic qubit can be distinguished by the photon counts Doherty et al. (2011, 2013). As the nuclear spin states are well isolated from the environment, the nuclear spin states can be read out by swapping the nuclear state to NV electronic states. Then the state can be detected using the above method Abobeih et al. (2022).
@@
Overall, the NV center systems have well-defined and controllable qubit states (Rule 1), and long coherence time (Rule4) and easy single-qubit measurement capabilities (Rule 5). However, the connectivity and two-NV entangling gates (Rule 2 and 3) are relatively weakly supported.
IV.2.3 Photonic qubits for quantum computing
@@
Optical photons are widely used in quantum communication and computing. However, their weak nonlinearity in optical nonlinear materials limits fast gate operations compared to microwave photons Byer (1974); Chen and Liu (1986), which benefit from Josephson junctions. Another distinct feature of optical photons used in quantum communication and quantum computing is that they are considered ‘flying qubits’, i.e., itinerant photons rather than stationary photons stored in an optical cavity mode. To avoid repeating the discussion covered in the previous sub-sections, we focus on itinerant photons used as qubits in quantum communication and quantum computing, especially in the MBQC model Raussendorf and Briegel (2001); Raussendorf et al. (2003); Browne and Rudolph (2005); Briegel et al. (2009), offering an alternative approach to supporting the AQM.
Qubit states.
@@
There are multiple ways of encoding quantum information into optical photons. For example, the quantum information can be encoded into the polarization of photons (polarization encoding) Turchette et al. (1995); Knill et al. (2001); Duan and Kimble (2004); O’Brien et al. (2003); Basset et al. (2021), the presence or absence of a photon (Fock encoding) Chuang and Yamamoto (1995), the presence of a photon in early or late time bins (time bin) Brendel et al. (1999); Jayakumar et al. (2014); Sun et al. (2022); Lee et al. (2019); Ortu et al. (2022), the presence of a photon in two modes with different frequencies (frequency encoding) Olislager et al. (2010); Lu et al. (2019); Sabattoli et al. (2022); Clementi et al. (2023), etc.
Operations.
@@
Universal control of a single qubit state varies based on encoding methods. Polarization-encoded photonic qubits can be controlled using passive optical elements like wave plates, beam splitters, phase shifters, and polarizers Kok et al. (2007). However, other encoding methods may require active elements and operations. For example, time-bin encoded qubits require delay lines and active optical switches to separate spatial modes, coherently convert photons between modes, and recombine them into time bins Kok et al. (2007).
@@
Supporting a universal gate set in photonic systems involves two approaches. One method is implementing an entangling gate, akin to other matter qubit systems. This gate can be executed probabilistically using photon measurements inducing nonlinearity Knill et al. (2001); O’Brien et al. (2003); Ralph et al. (2001); Zou et al. (2002); Pittman et al. (2002, 2003), or deterministically utilizing nonlinearity from strongly-coupled cavity-QED systems Duan and Kimble (2004); Kimble (2008); Zhan and Sun (2020). However, quantum gates suffer from imperfections. Probabilistic photon gates can achieve high fidelity through heralding but require more resources to implement Kok et al. (2007); O’Brien et al. (2003). Photonic gates using cavity-QED systems also face challenges due to imprecise control and photon absorption, making gate-based photonic QC difficult at the current stage.
@@
On the other hand, universal computing in photonic systems can be implemented by MBQC model Raussendorf and Briegel (2001); Raussendorf et al. (2003); Browne and Rudolph (2005); Briegel et al. (2009). In this model, universal computation is driven by measurements of selected qubits in a certain order on an existing entangled state. The entangled photonic states can be generated by time-delayed feedback Pichler et al. (2017); Sun et al. (2016); Zhan and Sun (2020); Shi and Waks (2021), or fusing small pieces of entangled states together Bartolucci et al. (2023), or by photon emission from entangled photonic emitters Lindner and Rudolph (2009); Economou et al. (2010); Buterakos et al. (2017); Russo et al. (2019).
@@
The physical support of a universal gate set can be disparate from other matter qubit systems discussed in the rest section. However, a universal gate used in the gate model, e.g., a CNOT gate, can be mapped to a sequence of measurements on a resource state Raussendorf and Briegel (2001); Raussendorf et al. (2003). Therefore, unlike gate-based models, after a gate operation, the quantum information is not carried by the same set of physical photonic qubits before the gate.
Connectivity.
@@
Long coherence time.
As photons do not interact with each other naturally, the photonic qubits can have long coherence times when they propagate in the vacuum. However, when they propagate in optical media, the material absorption causes photon loss, which is the leading error in photonic-based quantum systems. Another source of decoherence is the fluctuations of the optical paths, inducing phase noises to the photonic qubits. Stabilizing optical paths is essential in optical-based quantum systems Bernien et al. (2013); Pfaff et al. (2014); Hensen et al. (2015).
@@
Another way to mitigate the error is to enable error correction on photonic-based QC systems, which have been theoretically proposed and investigated Varnava et al. (2006, 2007); Buterakos et al. (2017); Zhan and Sun (2020); Hilaire et al. (2021); Zhan et al. (2023); Bartolucci et al. (2023). Generating tree graph states for photon loss error has been demonstrated in experiments Zhan and Sun (2020). However, due to the resource demand, it is still challenging to demonstrate the error correction functionality in photonic systems Li et al. (2015); Bartolucci et al. (2023).
Qubit readout.
Photon detection involves capturing photons by detectors, converting them into electric signals (e.g., current) Scully and Zubairy (1997). Utilizing photon detectors and linear optical devices like wave plates and polarizers, quantum information encoded in photonic qubits can be measured. For example, polarization-encoded photon qubits can be measured using polarizers followed by photon detectors, while the time-bin encoded photon qubits can be detected using optical switches to split them into two spatial modes, and then measured separately.
@@
IV.2.4 Quantum memory systems for photons
At the end of this section, we consider a type of physical system well-suited for buffering optical lights, which are used as optical quantum memories. These systems cannot support the AQM, however, still useful for QIP applications.
Qubit states.
There are different constructions of optical memories. Quantum memories using single atoms, or defect centers in solid-state systems can also be utilized as a computational qubit, as we discussed in Sec. IV.2.2. In this section, we examine two main types: atomic clouds and rare-earth-ion-doped crystals. Both types store quantum information in collective excitations of atoms or ions, rather than individual ones Lvovsky et al. (2009); Heshami et al. (2016); Fleischhauer et al. (2005); Lukin (2003); Ledingham et al. (2012); Ferguson et al. (2016); Alexander et al. (2006); Gorshkov et al. (2007).
@@ IV.2.4 Quantum memory systems for photons
Operations.
Realizing universal control of the Hilbert space spanned by the two states requires the collective control of all the atoms/ions inside the systems, which hinders the arbitrary transformations in the Hilbert space. However, by applying external optical light or electronic voltage to control the absorption of the atoms/ions, the incoming photons can be absorbed and converted into the collective excitation of the material, and then the absorbed photon can be re-emitted. Several techniques have been developed for these two operations, including electromagnetically-induced transparency (EIT) Fleischhauer et al. (2005); Lukin (2003); Hosseini et al. (2011); Dudin et al. (2013); Ding et al. (2013); Nicolas et al. (2014); Ding et al. (2015a, b); Parigi et al. (2015); Saunders et al. (2016); Katz and Firstenberg (2018); Hsiao et al. (2018); Jiang et al. (2019); Wang et al. (2019); Li et al. (2020); Dideriksen et al. (2021); Wang et al. (2022b); Meßner et al. (2023); Buser et al. (2022), controlled reversible inhomogeneous broadening (CRIB) Moiseev and Kröll (2001); Nilsson and Kröll (2005); Kraus et al. (2006); Alexander et al. (2006); Hétet et al. (2008), atomic frequency combs (AFC) McAuslan et al. (2012); Saglamyurek et al. (2011); Clausen et al. (2011); Ledingham et al. (2012); Zhou et al. (2012); Ferguson et al. (2016); Jin et al. (2022); Ma et al. (2021a); Seri et al. (2017); Kutluer et al. (2017); Laplane et al. (2017); Holzäpfel et al. (2020); Businger et al. (2020); Askarani et al. (2021); Ma et al. (2021b), and rephased amplified spontaneous emission (RASE) Ledingham et al. (2012, 2010).
Connectivity.
The multi-mode feature of the collective excitation used in atomic clouds and doped ions enables the integration of multiple memory cells into a single physical system Vernaz-Gris et al. (2018); Meßner et al. (2023). However, due to the control complexity, entangling different modes can be hardly implemented. Therefore, the connectivity of different memory modes is limited.
@@
Long coherence time.
The collective quantum state of the ensemble of atoms and ions can preserve the quantum information for a long time. For example, the $1/e$ decay lifetime of the stored light can reach $16$ s with dynamical decoupling in a cold atomic cloud using the EIT technique Dudin et al. (2013), while an AFC-based ion-doped crystal system an obtain storage lifetime $52.9$ min with dynamical decoupling Ma et al. (2021b).
Qubit readout.
@@
IV.3 Physical systems with co-design opportunities
@@
As the photonic system is challenging to perform two-qubit gates between physical qubits, Rule 2 and 3 of the AQM is not well supported between physical qubits. According to the unique support of the AQM by photonic qubit systems, MBQC is well-suited Bartolucci et al. (2023); Kok et al. (2007). In addition, although the entangling gates between photons are challenging, the photon hopping between coupled modes can be relatively easy to achieve, which provides the necessary connectivity (Rule 3) requirements for the application, such as quantum random walk models Tang et al. (2018) and sampling problems Zhong et al. (2020); Deng et al. (2023).
In contrast, NV center systems have long coherence times (Rule 4) and good support of the AQM Rule 1 and 5. Although the support of Rule 2 and 3 are limited, they are promising candidates for quantum memory and communication applications where a universal gate set isn’t necessary Bernien et al. (2013); Pfaff et al. (2014); Hensen et al. (2015); Borregaard et al. (2020); Pompili et al. (2021); Jing and Razavi (2022); Ruf et al. (2021); Liu et al. (2023b). Similar to the quantum memory systems for photons, where Rule 2 and 3 are not supported. In addition, the quantum memory systems for photons also have limited support to the readout capability (Rule 5). However, referring to Table. 2, they can be well suited to be used as quantum memory, which can not only be used in quantum communication and quantum networks Askarani et al. (2021); Jobez et al. (2016), but also in quantum memory of a heterogeneous QC architecture Liu et al. (2023b).
Superconducting, trapped ions, and neutral atom systems offer support of the complete AQM, making them promising candidates for not only universal quantum computing but also applications with stronger partial AQM demands on specific aspects. Compared to the superconducting qubit systems, trapped ions and neutral atoms have higher connectivity and longer coherence time, which also enable them to be used as quantum interconnect modules or quantum memory modules in future heterogenous QC systems with high fidelity coupling Stein et al. (2023a); Liu et al. (2023b). In addition, the high connectivity also makes them suitable for quantum annealing and quantum simulation applications Daley et al. (2022); Monroe et al. (2021); Ebadi et al. (2021).
V Conclusion
In conclusion, we refine the abstract qubit model, addressing the demands posed by quantum algorithms and applications, as well as the requirements on physical systems. However, achieving the ideal AQM is challenging for state-of-the-art physical systems. Instead, these systems can meet a less stringent set of requirements, utilizing techniques such as unitary decomposition, quantum error mitigation, and error correction methods, to compensate for the incomplete support of the AQM, albeit at the cost of more time and resources.

### v2

@@ 4.2.2 Nitrogen-Vacancy centers
Nitrogen-vacancy centers [138, 139, 140] combined with other solid-state defect centers have emerged as promising candidates for QIP. Defect color centers inside the solid state systems can be nicely fabricated and implanted inside the solid crystal, while they can have long coherence spin states that can be manipulated using microwave and optically readout.
a. Qubit states.
@@
The negatively charged NV centers have six electrons localized around the defect. The electronic ground state manifold consists of three spin-1 states. When there is no magnetic field applied, the state with $S_{1}=0$ and two states with $S_{z}=\pm 1$ splits by $2.87$ GHz [139, 140]. When the NV center electronic spin states are utilized to encode quantum information, a magnetic field is applied to further break the degeneracy of the two states with $S_{z}=\pm 1$, and use one of them with the $S_{z}=0$ state to form a qubit space [141, 142, 143]. Another way to encode quantum information is to encode into the nuclear spin states of the nearby atoms (Carbon or Nitrogen atoms) [144, 145, 146, 147, 148, 149].
b. Operations.
@@
The universal control of a qubit encoded into NV center electronic states can be achieved by coherent microwave drives [141, 142]. The universal control on the nuclear spin states can be achieved by addressing the nuclear spin states by manipulating the electronic states of the NV center via the hyperfine interaction [150, 151].
@@
However, implementing entangling gates between NV centers to support a full universal gate set for QC is challenging. Entanglement between two electronic spin qubits of NV centers can be entangled by photon heralded entanglement generation process [152, 153, 154, 146]. Entanglement gates between the electronic spin qubit and the nearby nuclear spin qubits can be performed by microwave drive on electronic spin degree of freedom [155, 156, 157, 158, 151].
@@
With the heralded entanglement between the electronic states of two NV centers, the entangled state can be used to enable gate operations between the nearby nuclear spin states. It can potentially be used to obtain a remote entangling gate for nuclear spin qubits, which can be used to build a quantum network [146].
c. Connectivity.
@@
Despite the possibility of heralded entanglement generation, two-qubit entanglement occurs relatively slowly, with a reported rate of $9$ Hz for remote entanglement between two NV centers [146]. Although improvements in photon emission, collection, and detection efficiency could enhance generation speed, the speed of generating entanglement of NV qubits remains slower compared to trapped ion and superconducting systems.
d. Long coherence time.
@@
The electronic qubit of NV centers can preserve a relatively long lifetime. Even at room temperature, the NV electronic states can preserve $1.8$ ms coherence time in an isotopically pure diamond crystal [159]. The nuclear spin states can have even longer lifetime [144, 145, 146, 147, 148, 149]. One challenge of NV center electronic states and nuclear spin states is the relatively short coherence time, which is due to the spin bath of the surrounding nuclear spins in the diamond crystal. However, the effect of the spin bath can be mitigated by dynamical decoupling [144, 160, 143, 151, 145]. With the help of dynamical decoupling, the coherence time of the electronic spin state can reach $1.58$ s [143], while the nuclear spin states can be extended to $63$ s [145, 151].
e. Qubit Readout.
@@
The measurement of NV center electronic states is implemented by optically pumping the states to the excited state manifold, and detecting the emitted photons. Due to the existence of a non-radiative relaxation path from the excited spin-00 states, the state of the NV electronic qubit can be distinguished by the photon counts [139, 140]. As the nuclear spin states are well isolated from the environment, nuclear spin states can be read out by swapping the nuclear state to NV electronic states. Then the state can be detected using the above method [151].
@@
Overall, the NV center systems have well-defined and controllable qubit states (Rule 1), long coherence time (Rule4), and easy single-qubit measurement capabilities (Rule 5). However, the connectivity and two-NV entangling gates (Rule 2 and 3) are relatively weakly supported.
4.2.3 Photonic qubits for quantum computing
@@
Optical photons are widely used in quantum communication and computing. However, their weak nonlinearity in optical nonlinear materials limits fast gate operations compared to microwave photons [161, 162], which benefit from Josephson junctions. Another distinct feature of optical photons used in quantum communication and quantum computing is that they are considered ‘flying qubits’, i.e., itinerant photons rather than stationary photons stored in an optical cavity mode. To avoid repeating the discussion covered in the previous sub-sections, we focus on itinerant photons used as qubits in quantum communication and quantum computing, especially in the MBQC model [14, 15, 16, 17], offering an alternative approach to supporting the AQM.
a. Qubit states.
@@
There are multiple ways of encoding quantum information into optical photons. For example, the quantum information can be encoded into the polarization of photons (polarization encoding) [163, 164, 165, 166, 167], the presence or absence of a photon (Fock encoding) [168], the presence of a photon in early or late time bins (time bin) [169, 170, 171, 172, 173], the presence of a photon in two modes with different frequencies (frequency encoding) [174, 175, 176, 177], etc.
b. Operations.
@@
Universal control of a single qubit state varies based on encoding methods. Polarization-encoded photonic qubits can be controlled using passive optical elements like wave plates, beam splitters, phase shifters, and polarizers [178]. However, other encoding methods may require active elements and operations. For example, time-bin encoded qubits require delay lines and active optical switches to separate spatial modes, coherently convert photons between modes, and recombine them into time bins [178].
@@
Supporting a universal gate set in photonic systems involves two approaches. One method is implementing an entangling gate, akin to other matter qubit systems. This gate can be executed probabilistically using photon measurements inducing nonlinearity [164, 166, 179, 180, 181, 182], or deterministically utilizing nonlinearity from strongly-coupled cavity-QED systems [165, 183, 184]. However, quantum gates suffer from imperfections. Probabilistic photon gates can achieve high fidelity through heralding but require more resources to implement [178, 166]. Photonic gates using cavity-QED systems also face challenges due to imprecise control and photon absorption, making gate-based photonic QC difficult at the current stage.
@@
On the other hand, universal computing in photonic systems can be implemented by MBQC model [14, 15, 16, 17]. In this model, universal computation is driven by measurements of selected qubits in a certain order on an existing entangled state. The entangled photonic states can be generated by time-delayed feedback [185, 186, 184, 187], or fusing small pieces of entangled states together [188], or by photon emission from entangled photonic emitters [189, 190, 191, 192].
@@
The physical support of a universal gate set can be disparate from other matter qubit systems discussed in the rest section. However, a universal gate used in the gate model, e.g., a CNOT gate, can be mapped to a sequence of measurements on a resource state [14, 15]. Therefore, unlike gate-based models, after a gate operation, the quantum information is not carried by the same set of physical photonic qubits before the gate.
c. Connectivity.
@@
d. Long coherence time.
As photons do not interact with each other naturally, the photonic qubits can have long coherence times when they propagate in the vacuum. However, when they propagate in optical media, the material absorption causes photon loss, which is the leading error in photonic-based quantum systems. Another source of decoherence is the fluctuations of the optical paths, inducing phase noises to the photonic qubits. Stabilizing optical paths is essential in optical-based quantum systems [152, 153, 154].
@@
Another way to mitigate the error is to enable error correction on photonic-based QC systems, which have been theoretically proposed and investigated [193, 194, 191, 184, 195, 196, 188]. Generating tree graph states for photon loss error has been demonstrated in experiments [184]. However, due to the resource demand, it is still challenging to demonstrate the error correction functionality in photonic systems [197, 188].
e. Qubit readout.
Photon detection involves capturing photons by detectors, converting them into electric signals (e.g., current) [198]. Utilizing photon detectors and linear optical devices like wave plates and polarizers, quantum information encoded in photonic qubits can be measured. For example, polarization-encoded photon qubits can be measured using polarizers followed by photon detectors, while the time-bin encoded photon qubits can be detected using optical switches to split them into two spatial modes, and then measured separately.
@@
4.2.4 Quantum memory systems for photons
At the end of this section, we consider a type of physical system well-suited for buffering optical light, which is used as optical quantum memories. These systems cannot support the AQM, however, still useful for QIP applications.
a. Qubit states.
There are different constructions of optical memories. Quantum memories using single atoms, or defect centers in solid-state systems can also be utilized as a computational qubit, as we discussed in Sec. 4.2.2. In this section, we examine two main types: atomic clouds and rare-earth-ion-doped crystals. Both types store quantum information in collective excitations of atoms or ions, rather than individual ones [199, 200, 201, 202, 203, 204, 205, 206].
@@ 4.2.4 Quantum memory systems for photons
b. Operations.
Realizing universal control of the Hilbert space spanned by the two states requires the collective control of all the atoms/ions inside the systems, which hinders the arbitrary transformations in the Hilbert space. However, by applying external optical light or electronic voltage to control the absorption of the atoms/ions, the incoming photons can be absorbed and converted into the collective excitation of the material, and then the absorbed photon can be re-emitted. Several techniques have been developed for these two operations, including electromagnetically-induced transparency (EIT) [201, 202, 207, 208, 209, 210, 211, 212, 213, 214, 215, 216, 217, 218, 219, 220, 221, 222, 223], controlled reversible inhomogeneous broadening (CRIB) [224, 225, 226, 205, 227], atomic frequency combs (AFC) [228, 229, 230, 203, 231, 204, 232, 233, 234, 235, 236, 237, 238, 239, 240], and rephased amplified spontaneous emission (RASE) [203, 241].
c. Connectivity.
The multi-mode feature of the collective excitation used in atomic clouds and doped ions enables the integration of multiple memory cells into a single physical system [242, 222]. However, due to the control complexity, entangling different modes can be hardly implemented. Therefore, the connectivity of different memory modes is limited.
@@
d. Long coherence time.
The collective quantum state of the ensemble of atoms and ions can preserve the quantum information for a long time. For example, the $1/e$ decay lifetime of the stored light can reach $16$ s with dynamical decoupling in a cold atomic cloud using the EIT technique [208], while an AFC-based ion-doped crystal system an obtain storage lifetime $52.9$ min with dynamical decoupling [240].
e. Qubit readout.
@@
4.3 Physical systems with co-design opportunities
@@
As the photonic system is challenging to perform two-qubit gates between physical qubits, Rule 2 and 3 of the AQM is not well supported between physical qubits. According to the unique support of the AQM by photonic qubit systems, MBQC is well-suited [188, 178]. In addition, although the entangling gates between photons are challenging, the photon hopping between coupled modes can be relatively easy to achieve, which provides the necessary connectivity (Rule 3) requirements for the application, such as quantum random walk models [60] and sampling problems [243, 244].
In contrast, NV center systems have long coherence times (Rule 4) and good support of the AQM Rule 1 and 5. Although the support of Rule 2 and 3 are limited, they are promising candidates for quantum memory and communication applications where a universal gate set isn’t necessary [152, 153, 154, 245, 146, 246, 247, 51]. Similar to the quantum memory systems for photons, where Rule 2 and 3 are not supported. In addition, the quantum memory systems for photons also have limited support to the readout capability (Rule 5). However, referring to Table. 2, they can be well suited to be used as quantum memory, which can not only be used in quantum communication and quantum networks [239, 248], but also in quantum memory of a heterogeneous QC architecture [51].
Superconducting, trapped ions, and neutral atom systems offer support for the complete AQM, making them promising candidates for not only universal quantum computing but also applications with stronger partial AQM demands on specific aspects. Compared to the superconducting qubit systems, trapped ions and neutral atoms have higher connectivity and longer coherence time, which also enable them to be used as quantum interconnect modules or quantum memory modules in future heterogenous QC systems with high fidelity coupling [49, 51]. In addition, the high connectivity also makes them suitable for quantum annealing and quantum simulation applications [68, 65, 67].
5 Conclusion
In conclusion, we refine the abstract qubit model, addressing the demands posed by quantum algorithms and applications, as well as the requirements of physical systems. However, achieving the ideal AQM is challenging for state-of-the-art physical systems. Instead, these systems can meet a less stringent set of requirements, utilizing techniques such as unitary decomposition, quantum error mitigation, and error correction methods, to compensate for the partial support of the AQM, albeit at the cost of more time and resources.